OpenStudy (anonymous):

convert x^2 -x -8 into standard form

OpenStudy (anonymous):

yes. I don't know how to convert it to a(x-h)^2

OpenStudy (anonymous):

u need the start form of a polynomial?

OpenStudy (anonymous):

yes please!1

OpenStudy (anonymous):

it's already in start form :D

OpenStudy (anonymous):

sry standard :D

OpenStudy (anonymous):

i mean the other form

OpenStudy (anonymous):

we don't have another form...you wanted the standard form and \(x^2 -x -8\) is standard form itself.

OpenStudy (anonymous):

like y=a(x-h)^2+k

OpenStudy (anonymous):

i havn't seen that form and i don't think that it can be converted to that form...

OpenStudy (anonymous):

are you sure??

OpenStudy (anonymous):

Express the quadratic function f(x) = x2 − x − 8 in standard form. f(x) =

OpenStudy (anonymous):

i'm pretty sure...i havn't seen that form.

OpenStudy (the_fizicx99):

That's not standard form. y = a(x - h)^2 + k is vertex form

OpenStudy (the_fizicx99):

Do you want me to explain it to you?

OpenStudy (anonymous):

@tHe_FiZiCx99 yes please i'm sorry I took so long